### v1

# The effect of control barrier functions on energy transfers in controlled physical systems

Abstract
Using a port-Hamiltonian formalism, we show the qualitative and quantitative effect of safety-critical control implemented with control barrier functions (CBFs) on the power balance of controlled physical systems. The presented results will provide novel tools to design CBFs inducing desired energetic behaviors of the closed-loop system, including nontrivial damping injection effects and non-passive control actions, effectively injecting energy in the system in a controlled manner. Simulations validate the stated results.
Index Terms:
Control barrier functions, port-Hamiltonian systems, energy-based control, nonlinear control
I Introduction
Control objectives in engineering applications are becoming increasingly complex in terms of their semantic and mathematical description. An instructive example to think about is safe collaboration between humans and robots [Robla-Gomez2017WorkingEnvironments, Hamad2023AInteraction, Siciliano2016SpringerRobotics], where control goals are often represented by multi-objective functions representing the specific task to be executed. These objectives (which are sometimes unknown and implicitly designed through e.g. control by human demonstrations) are non-trivially intertwined with safety specifications, aimed at representing critical hazards whose prevention needs to be certified before any tentative of real-world application. These safety specifications are often conflicting with the main task to be executed, and the semantic distinction between the “main task”, represented by objective functions, and the “safety specifications”, represented by constraints on the state of the system in the optimization problem, is ubiquitous in the control-theoretic literature [Ames2019ControlApplications, Wabersich2023Data-DrivenSystems, Ferraguti2022SafetyApproach]. This view of safety-critical control applications has led to the concept of safety filters [Wabersich2023Data-DrivenSystems], which are computational units with the goal of transforming a nominal control input (designed for the main task, disregarding the safety requirements) into a new control input which i) minimally modifies the nominal controller and ii) ensures the safety requirements. One of the most popular techniques for implementing these safety filters is represented by control barrier functions (CBF) [Ames2019ControlApplications, Ames2017ControlSystems], an algorithm which aims to modify a nominal control input to achieve forward invariance of a safe set.
In this paper, we perform an in-depth analysis of the effect of control barrier functions (CBFs) on energy transfers in physical systems, inducing novel control design tools aimed at regulating the energy transfers along the closed-loop system. The idea that an energetic description captures the behavioral property of a system is present in some branches of control theory, especially in the so-called energy-based or energy-aware control literature [Duindam2009ModelingSystems, Stramigioli2015Energy-AwareRobotics, Califano2022OnSystems, Ortega2001PuttingControl]: if we are able to control the energy transfers undergoing the controlled system, then we can achieve desired behaviors, going beyond mere stabilization purposes. This aspect is particularly valid in collaborative robotic applications that take place in unstructured environments, possibly in the presence of humans, since energy transfers between the robot and its surroundings encode information on performance, safety and energy efficiency at the same time [Tadele2014CombiningRobots, Laffranchi2009SafeControl, Zacharaki2020SafetySurvey, Califano2022OnSystems, Tadele2014ThePublications]. This narrative of “putting energy back in control” [Ortega2001PuttingControl] has been the driving motivation in the development of control schemes in the port-Hamiltonian formalism [Duindam2009ModelingSystems].
Port-Hamiltonian (pH) systems are control-affine systems that serve best to represent autonomous and controlled physical systems, since their structural properties directly relate to their underlying energetic structures. PH systems basically encompass all physical systems of interest, finite- or infinite-dimensional, and from any physical domain [Duindam2009ModelingSystems, Rashad2020TwentyReview]. As a consequence, the use of this formulation does not introduce any restriction to the class of physical systems that can be modeled, while at the same time it introduces technical and intuitive advantages when studying the effects of CBF algorithms on the energy balance of the system. Among the possible energy-aware control strategies that can be developed with the tools presented in this work, we will present novel damping injection schemes, in which the CBF algorithm induces passive control actions that extract energy from the controlled system. Furthermore, we introduce new schemes which effectively inject energy in the controlled system, going beyond passive designs and giving new perspectives to achieve desired closed-loop behaviors.
In Sec. II the background on pH systems and CBFs is reviewed. In Sec. III the main contribution of this letter is given and applied to mechanical systems in Sec. IV. In Sec. V we show numerical simulations and Sec. VI contains conclusions and future work.
II Background
We refer to [vanderSchaftL2, Duindam2009ModelingSystems] regarding pH systems and to [Ames2019ControlApplications, Xu2015RobustnessControl, Ames2017ControlSystems, Singletary2021Safety-CriticalSystems] regarding CBF for references that fully cover the presented background. The specialisation of CBFs to pH systems is a contribution on its own present in this section.
II-A Port-Hamiltonian systems
The input–state–output representation of a port–Hamiltonian (pH) system is:
$\begin{cases}\dot{x}=(J(x)-R(x))\partial_{x}H(x)+g(x)u\\ y=g(x)^{\top}\partial_{x}H(x)\end{cases}$ (1)
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the input, $g(x)$ is the input matrix, $J(x)=-J(x)^{\top}$ and $R(x)=R(x)^{\top}\geq 0$ are respectively skew-symmetric and positive semi-definite symmetric matrices representing the power-preserving and the dissipative components of the system. The non-negative function $H:\mathcal{D}\to\mathbb{R}^{+}$ is called the Hamiltonian and maps the state into the total physical energy of the system. $\partial_{x}H(x)\in\mathbb{R}^{n}$ denotes the gradient of $H$, represented as a column, and $\partial_{x}^{\top}H(x)$ denotes its transposed. When not explicitly stated, all described variables are assumed to have a degree of continuity such that the right-hand side of (1) is locally Lipschitz, to guarantee the existence and uniqueness of the solutions.
We now present geometric properties of port-Hamiltonian systems that are relevant in this paper. The use of coordinate-free notation will induce intuitive and technical advantages in the interpretability of the expressions and in the calculations.
II-A1 Power-preserving structure
The skew-symmetric matrix operator $J(x)$ induces the bracket {,}$\{\cdot,\cdot\}_{J}$, a skew-symmetric bilinear map that takes as input two smooth scalar functions on the state space to produce another scalar function. Given two scalar functions $A:\mathcal{D}\to\mathbb{R}$ and $B:\mathcal{D}\to\mathbb{R}$, the bracket is defined as $\{A,B\}_{J}:=\partial_{x}^{\top}A(x)J(x)\partial_{x}B(x)$. When the second slot of the bracket is fed with the Hamiltonian $H$, the bracket completely represents the conservative Hamiltonian dynamics obtained from (1) setting $R(x)=0$ and $u=0$. In this case, for any function $A:\mathcal{D}\to\mathbb{R}$, its variation along the solution of (1) is $\dot{A}=\{A,H\}_{J}$, and as a particular case the conservation of energy is encoded in the skew-symmetry of the bracket since $\dot{H}=\{H,H\}_{J}=0$.
II-A2 Dissipation structure
We denote by the bracket $[\cdot,\cdot]_{Y}$ the bilinear map $[A,B]_{Y}:=\partial_{x}^{\top}A(x)Y(x)\partial_{x}B(x)$ for a state-dependent symmetric matrix $Y(x)\in\mathbb{R}^{n\times n}$. For system (1), when $Y(x)=R(x)$ and the bracket is fed twice with the Hamiltonian, this bracket represents dissipated power due to the dissipative effects modeled in $R(x)$, i.e., $[H,H]_{R}=\partial_{x}^{\top}H(x)R(x)\partial_{x}H(x)\geq 0$.
II-A3 Passivity
The input $u$ and the output $y$ in (1) are co-located, in the sense that the variation of energy due to the input is $\partial_{x}^{\top}H(x)g(x)u=y^{\top}u$. It follows that system (1) is passive with the Hamiltonian $H(x)$ as storage function and input-output pair $(u,y)$:
$\dot{H}=\partial_{x}^{\top}H(x)\dot{x}=\underbrace{\{H,H\}_{J}}_{0}-\underbrace{[H,H]_{R}}_{\geq 0}+y^{\top}u\leq y^{\top}u.$ (2)
The latter power balance is a statement of energy conservation for the physical system (1), where it should be appreciated how the pH structure displays all the possible ways in which physical energy can flow along the system: the skew-symmetric structure $J$ represents pure routing of energy, the dissipative structure $R$ dissipated energy, and the duality product $y^{\top}u$ represents the instantaneous power injected by the input. Passivity condition (2) states that the variation of energy in the system is bounded by the injected power.
II-A4 Control
From a system-theoretic perspective passivity implies stability under weak conditions: $H(x)$ is a valid Lyapunov candidate since $u=0$ implies $\dot{H}\leq 0$. Since (1) is time-invariant, the autonomous system converges to the largest invariant subset of $\{x\in\mathcal{D}|[H,H]_{R}=0\}$ which generally depends on the dissipation matrix $R(x)$ and the Hamiltonian $H(x)$. Many pH inspired control schemes have been proposed, with the rationale of obtaining a closed-loop system in the form (1) with Hamiltonian and system matrices corresponding to desired behaviors.
II-B Control barrier functions
Control barrier functions (CBFs) represent a technique to guarantee forward invariance of a set $\mathcal{C}$, normally called safe set. We present the standard background for CBFs directly specialised to port-Hamiltonian systems (1), providing interpretability of CBF-based algorithms from an energetic perspective.
The goal is to design a state feedback $u=k(x)$ resulting in the closed-loop system $\dot{x}=(J(x)-R(x))\partial_{x}H+g(x)k(x)$ such that
$\forall x(0)\in\mathcal{C}\implies x(t)\in\mathcal{C}\,\,\,\forall t>0.$ (3)
The safe set $\mathcal{C}$ is built as the superlevel set of a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$, i.e., $\mathcal{C}=\{x\in\mathcal{D}:h(x)\geq 0\}.$ The function $h(x)$ is then defined as a CBF on $\mathcal{D}$ if $\partial_{x}h(x)\neq 0,\forall x\in\partial\mathcal{C}$ and
$\sup_{u\in\mathcal{U}}[\dot{h}(x,u)]\geq-\alpha(h(x))$ (4)
for all $x\in\mathcal{D}$ and some extended class $\mathcal{K}$ function [^1] A function $\alpha:(-b,a)\to(-\infty,\infty)$ with $a,b>0$, which is continuous, strictly increasing, and $\alpha(0)=0$. $\alpha$. Here we used $\dot{h}=\partial_{x}^{\top}h(x)\dot{x}$ to denote the time derivative of $h$ along the solution of (1)), which results in:
$\dot{h}(x,u)=\{h,H\}_{J}-[h,H]_{R}+\partial_{x}^{\top}h(x)g(x)u.$ (5)
The link between the existence of a CBF and the forward invariance of the related safe set is established by the following key result.
Theorem 1 ([Ames2017ControlSystems], here applied to pH systems (1)).
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1). Any locally Lipschitz controller $u=k(x)$ such that $\{h,H\}_{J}-[h,H]_{R}+\partial_{x}^{\top}h(x)g(x)k(x)\geq-\alpha(h(x))$ provides forward invariance of the safe set $\mathcal{C}$. Additionally the set $\mathcal{C}$ is asymptotically stable on $\mathcal{D}$.
The way controller synthesis induced by CBFs are implemented is to use them as safety filters, transforming a nominal state-feedback control input $u_{\textrm{nom}}(x)$ into a new state-feedback control input $u^{*}(x)$ in a minimally invasive fashion in order to guarantee forward invariance of $\mathcal{C}$. In practice, the following Quadratic Program (QP) is solved:
$u^{*}(x)=$ $\operatorname*{argmin}_{u\in\mathcal{U}}\quad||u-u_{\textrm{nom}}(x)||^{2}$ (6)
[…]
The transformation of the desired control input $u_{\textrm{nom}}(x)$ in $u^{*}(x)$ by solving (6) is denoted as safety-critical control. A last result that will be crucially used in this work is the following lemma.
Theorem 2 ([Xu2015RobustnessControl, Singletary2021Safety-CriticalSystems], here applied to pH systems (1)).
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1) and assume $\mathcal{U}=\mathbb{R}^{m}$. Define $\Psi(x):=\dot{h}(x,u_{\textup{nom}}(x))+\alpha(h(x))$. Omitting functional dependencies for readability, a closed-form solution for (6) is given by $u^{*}(x)=u_{\textup{nom}}(x)+u_{\textup{safe}}(x)$, where
$u_{\textup{safe}}(x)=\begin{cases}\begin{aligned} &-\frac{g^{\top}\partial_{x}h}{[h,h]_{gg^{\top}}}\Psi&\textup{if}\,\Psi<0\\ &\qquad\qquad 0&\textup{if}\,\Psi\geq 0\end{aligned}\end{cases}.$ (7)
For the remainder of this paper we will assume $u_{\textrm{nom}}=0$. This choice is consistente with the way controllers are designed in the pH framework. In fact, as reported in Sec. II, the form (1) can already represent controlled physical systems performing some nominal task with a desired energetic behavior.
III Putting energy back in control through CBFs
In this section, we discuss how CBF-induced safety-critical control influences the power balance of the controlled physical system. We start presenting a closed-form solution for the power injected by the safety-critical controller for controlled pH systems.
Theorem 3.
Let the assumptions of Theorem 2 hold. The power balance for the pH system (1), controlled with input $u_{\textrm{nom}}(x)$ undergoing safety-critical control (6) induced by a CBF $h(x)$, is
$\dot{H}=-[H,H]_{R}-\mathbbm{1}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi,$ (8)
where $\Psi=\{h,H\}_{J}-[h,H]_{R}+\alpha(h(x))$ and the indicator function $\mathbbm{1}_{\Psi<0}$ returns $1$ if $\Psi<0$ ($u_{\textrm{safe}}\neq 0$) and $1$ otherwise ($u_{\textrm{safe}}=0$).
Proof.
The results follows from an explicit calculation of the power balance (2) using $u=u_{\textup{safe}}(x)$ given by (7). ∎
We indicate the instantaneous power injection induced by the safety-critical controller (the last addend in (8)) with
$P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi$ (9)
where the notation stresses the fact that this term depends both on the CBF $h$ and on the function $\alpha$, a parameter of the CBF algorithm whose role will be understood in energetic terms in the sequel.
III-A Preserving stability: CBF-based damping injection
As reported in Sec. II-A4, if the Hamiltonian $H$ qualifies as a Lyapunov function, the minimum of $H$ is a stable equilibrium (the time derivative of $H$ is given by (8) with $P_{h,\alpha}=0$). Natural questions that can be answered are: under what conditions does the safety-critical control preserve the stability of the underlying controlled system? If stability is preserved, what is the effect of safety-critical control on the rate of convergence to the equilibrium? As (8) holds true, in case $H$ would still qualify as a Lyapunov function after the safety-critical control, the closed-loop system would converge to the largest invariant subset in $\{x\in\mathcal{D}|P_{h,\alpha}-[H,H]_{R}=0\}$. The following result gives a useful sufficient condition for preservation of local stability.
Proposition 1.
Suppose $H$ is a Lyapunov function for the pH system (1). Suppose that a safety-critical control algorithm induced by a CBF $h(x)$ acts on the system under the assumptions of Theorem 2. The stability of the original system is preserved if $\mathbbm{1}_{\Psi<0}[H,h]_{gg^{\top}}\leq 0$ in a neighborhood of the equilibrium.
Proof.
The result follows by working out analitically the sufficient condition under which $P_{h,\alpha}$, and as such the whole right-hand side of (8), is non-positive. ∎
The condition $P_{h,\alpha}\leq 0$ forces the safety-critical controller to “act as a damper”, possibly helping dissipative dynamics encoded in (1) in the stabilization task, normally referred to as damping injection in the pH framework. The closed-form expression (9) can be used to design and interpret the non-trivial damping effects that specific CBFs have on the controlled system.
III-B Beyond damping injection
In a safety-critical framework, we find a strong conceptual distinction between the “main task” to be executed and the “safety constraints” which have to be guaranteed in any circumstance. Although this narrative is useful in many applications and is widespread in the control-theoretic literature, the main controller and the CBF-induced term must not necessarily have antagonistic roles, as seen e.g., in previous section where the safety-critical controller can act as a smart damping injection mechanism. More in general, the power term induced by CBFs (9) induces a constructive way to “put energy back in control”, in the sense envisioned in energy-based methods (see e.g., [Ortega2001PuttingControl] and [Stramigioli2015Energy-AwareRobotics]), also going beyond purely passive designs and possibly injecting positive power in the system if the task requires it. To show the potential of the approach in the context of pH systems, it is instructive to design a CBF that encodes as a safe set the fact that the system has to stay below or above a certain energy limit $c$, that is, $h(x)=\pm(c-H(x))$. The following example shows what a control designer can gain in terms of interpretability of the safety-critical control action, as well as the role of the class $\mathcal{K}$ function $\alpha$ in the algorithm.
Example 1 (Limiting total energy from above and below).
Consider the case $R=0$, that is, the system (1) evolves on energetic isolines.
Using a CBF $h(x)=c-H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\leq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(c-H(x)).$ (10)
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$, while if $x\notin\mathcal{C}$ the safety-critical control extracts the negative power $P_{h}=\alpha(h)$, which can be modulated with $\alpha$. In case the $\alpha$ is linear, i.e., $\alpha(h)=\gamma h,\,\gamma\in\mathbb{R}^{+}$, the constant $\gamma$ acts as a proportional gain on the “energy error” $c-H(x)$ measuring the distance from the closest energy value that the system is ultmately allowed to stay in. The (simple) stability of the minimum of the Hamiltonian is clearly preserved as $[h,H]_{gg^{\top}}=-[H,H]_{gg^{\top}}\leq 0$.
Using instead $h(x)=-c+H(x)$, i.e., $\mathcal{C}=\{x\in\mathcal{D}|H(x)\geq c\}$, the power equality (8) becomes:
$\dot{H}=P_{h,\alpha}=\mathbbm{1}_{\Psi<0}(\alpha(c-H(x)),\qquad\Psi=\alpha(H(x)-c).$ (11)
If $x\in\mathcal{C}$ then $P_{h,\alpha}=0$, but now if $x\notin\mathcal{C}$, $P_{h,\alpha}$ is positive, corresponding to a non-passive action, as the safety-critical controller needs to inject power into the system to reach the set $\mathcal{C}$, which can be again modulated by $\alpha$.
IV Mechanical systems and energy-based CBFs
In order to better comprehend the proposed methodology and discuss how it complements standard energy-based approaches, let us specialise system (1) to fully actuated mechanical systems, where we introduce the state $x=(q^{\top},p^{\top})^{\top}\in\mathbb{R}^{2n}$ as canonical Hamiltonian coordinates on the cotangent bundle of the configuration manifold of the system. In particular $q\in\mathbb{R}^{n}$ is the vector of generalized coordinates. $p\in\mathbb{R}^{n}$ denotes the generalized momenta, $p:=M(q)\dot{q}$, where $M(q)=M(q)^{\top}>0$ is the positive definite inertia matrix of the system. The equations of motions in canonical form are given by (1) with
$J(x)-R(x)=\begin{bmatrix}0&I_{n}\\ -I_{n}&-D\end{bmatrix},~{}g(x)=\begin{bmatrix}0\\ B\end{bmatrix}$
resulting in
[…]
where $H:\mathbb{R}^{2n}\to\mathbb{R}$ is the total energy (Hamiltonian)
$H(q,p)=K_{e}(q,p)+V(q),$
where $K_{e}(q,p)=\frac{1}{2}p^{\top}M^{-1}(q)p$ is the kinetic energy, $V:\mathbb{R}^{n}\to\mathbb{R}$ maps the position state to conservative potentials (gravity, elastic effects), $D=D^{\top}\geq 0$ takes into account dissipative and friction effects, $B\in\mathbb{R}^{n\times n}$ is the (full rank) input matrix [^2] For lighten notation we hide possible state dependencies on $D$ and $B$., $I_{n}$ indicates the $n\times n$ identity matrix and non specified dimensions of matrices, comprising those with only zero entries and denoted with the symbol “00”, are clear from the context.
A safety-critical control action with CBF $h(q,p)$ specialises then the total power balance (8) into
$\dot{H}=-\dot{q}^{\top}D{\dot{q}}+P_{h,\alpha},$ (13)
where
$P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}\frac{\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}}{\frac{\partial^{\top}h}{\partial p}BB^{\top}\frac{\partial h}{\partial p}}\Psi,$ (14)
and
$\Psi=\{h,H\}-\frac{\partial^{\top}h}{\partial p}D{\dot{q}}+\alpha(h(q,p)),$ (15)
[…]
The previous expressions allow to design safety-critical controllers with a qualitiative and quantitative insight on the effect of a certain CBF $h(q,p)$ on the power balance of the controlled mechanical system. An interesting class of CBFs which deserves some discussion is
$h(q,p)=\pm K_{e}(q,p)+\bar{h}(q)+c,$ (16)
where $\bar{h}$ is a smooth function of the configuration variable only and $c$ is a real number.
The sufficient condition of Proposition 1 on stability-preserving safety-critical control (i.e., the condition on which safety-critical control always injects damping in the controlled mechanical system) reduces to
$\mathbbm{1}_{\Psi<0}\bigg{(}\dot{q}^{\top}BB^{\top}\frac{\partial h}{\partial p}\bigg{)}\leq 0$ (17)
Since $\partial K_{e}(q,p)/\partial p=\dot{q}$, a subclass of CBFs in the form (16) which satisfy this condition is:
$h(q,p)=-K_{e}(q,p)+\bar{h}(q)+c,$ (18)
and the amount of instantaneous dissipated power results in:
$P_{h,\alpha}=\mathbbm{1}_{\Psi<0}\Psi\leq 0.$ (19)
Remark 1.
CBFs (18) encompass those referred to as energy-based CBFs in [Singletary2021Safety-CriticalSystems], defined as (18) with $c=0$ and $\bar{h}=\alpha_{E}h_{k}(q)$ where $h_{k}(q)$ is a function whose superlevel safe set represents purely kinematic constraints (e.g., obstacle avoidance). In [10136379] the safety-critical control induced by CBFs in the form (18) was proven to preserve passivity of nominal controllers, similarly to the discussed in Proposition 1, where $H$ is assumed to be a Lyapunov function.
Finally, we state the following insightful result.
Proposition 2.
Given a mechanical pH system (12) undergoing a safety-critical control induced by a CBF in the form (16), the expression for $\Psi$ in (15) specialises to
$\Psi=\mp\dot{V}+\dot{\bar{h}}+\dot{q}^{\top}D{\dot{q}}+\alpha(h(q,p))$ (20)
Proof.
We use properties of skew-symmetry and bilinearity of the Poisson bracket to calculate the expression $\{h,H\}$ for (16): {±$\{\pm K_{e}+\bar{h},K_{e}+V\}=\{\pm K_{e},V\}+\{-K_{e},\bar{h}\}=\{\pm K_{e},V\mp\bar{h}\}$. Using the definition of the bracket and the fact that $V$ and $\bar{h}$ do not depend on $p$, one gets {±$\{\pm K_{e},V\mp\bar{h}\}=\mp\frac{\partial^{\top}K_{e}}{\partial p}\frac{\partial(V\mp\bar{h})}{\partial q}=\mp\frac{\partial^{\top}(V\mp\bar{h})}{\partial q}\dot{q}=\dot{\bar{h}}\mp\dot{V}.$
This result provides further interpretability in the structure of $\Psi$, whose negativity determines when the safety-critical control starts to act, and whose absolute value determines the amount of power which is injected (if (16) is taken with the plus) or dissipated (if (16) is taken with the minus) instantaneously in the system. We give practical insight with the following examples.
Example 2 (Example 1 applied to mechanical systems).
Notice that for CBFs of the form (18) (pure damping injection), the first case in Example 1 is retrieved if we use $\bar{h}=-V$, and thus $\{h,H\}=0$ and the value of $\Psi$ reduces to (10) if friction is ignored. Similarly, one can retrieve the second case in Example 1, i.e., choosing a CBF (16) with the plus sign and $\bar{h}=V$.
The latter example is a particular case in which the positivity of $\Psi$ depends only on whether the system is in the safe set, but more in general the result in Proposition 2 allows one to get information about the way safety-critical control acts in terms of energy transfer and use that information in the CBF design. As an example, we present a case of safety-critical control that involves the prevention of several physical safety hazards in robotics by limiting the total kinetic energy of the system.
Example 3 (Limiting kinetic energy).
If one wants to limit kinetic energy in a mechanical system (12), the candidate CBF is $h(q,p)=-K_{e}(q,p)+c$, a particular case of (18) with $\bar{h}=0$. Notice that in this case, if there is no friction, we obtain $\Psi=\dot{V}+\alpha(h(q,p))$, which encodes the information on when the safety-critical effects occur, how much damping is injected in the system, and how the design of the class $\mathcal{K}$ function $\alpha$ could be used to influence the resulting dynamics.
V Simulations
Figure 1: Mass-Spring system. Bounding energy from above (left) and below (right). From above: i) phase space trajectory, ii) CBF, iii) safety-critical control input, and iv) total energy.
Figure 2: Mass-Spring system. Limiting kinetic energy starting from inside (left) and outside (right) the safe set. From above: i) phase space trajectory, ii) CBF, iii) safety-critical control input, and iv) total energy.
Figure 3: Double pendulum. Limiting kinetic energy starting from inside (left) and outside (right) the safe set. From above: i) CBF, ii) safety-critical control input, and iii) total energy.
In this section, we display and discuss numerical simulations of mechanical systems undergoing safety-criticl control with CBFs in the form (16), previously discussed in Examples 2 and 3. When the phase space is plotted, the safe set is represented in white and its complement is represented in gray. The class $\mathcal{K}$ function $\alpha$ is chosen linearly, i.e., $\alpha(h)=\gamma h$, and simulations are performed for different positive scalars $\gamma$.
V-1 Bounding total energy from below and above
To display the validity and the potential of the proposed results we initially consider as a plant a simple oscillator (mass-spring) with no dissipation, i.e., system (12) with $D=0$, $B=1$, $H=\frac{p^{2}}{2m}+\frac{kq^{2}}{2}$, $k=0.5$ and $m=2$. The results are shown in Fig. 1. The left column implements a safety-critical controller given by a CBF bounding the total energy from above ($h=-H+c$) while the second column uses a CBF bounding the total energy from below ($h=H-c$), with energy limit $c=10$. It is interesting to see how the proposed approach generates new ways to implement mechanical oscillations at the desired energy level. As theoretically derived from example 1, we see that the safety critical input is triggered once the system is outside the safe set, and we see how the $\gamma$ parameter can be set to regulate the amount of power the controller dissipates or injects.
V-2 Bounding kinetic energy
Fig. 2 shows simulations with the same plant as before, but with a CBF designed to limit only the kinetic energy, i.e., $h=-K_{e}+c$ where $c=20$, as explained in Example 3. The left column corresponds to experiments in which the system starts in the safe set, and the right column to one in which the system starts outside the safe set. These simulations clearly show both the invariance and the asymptotic stability of the safe set. Looking at the total energy, it appears clearly that, as predicted by the theory, the safety-critical controller acts purely as a damper, with the amount of damping which be regulated with $\gamma$. As explained in Example 3, the activation of safety-critical effects is no longer trivially related to whether the system is in the safe set or not, and can be inspected by checking the positivity of the constraint (20).
Finally, in Fig. 3 we show simulations for a double pendulum with masses $m_{1}=m_{2}=2$ and lengths $l_{1}=l_{2}=1$, without natural dissipation, in a gravitational field $V(q)$ (the zero potential corresponds to the reference configuration in which the pendulum is in the horizontal position). This system is described by (12) with $D=0$, $B=I_{2}$ and $H=\frac{p^{\top}M^{-1}(q)p}{2}+V(q)$. The CBF is designed to limit only the kinetic energy, i.e., $h=-K_{e}+c$ where $c=10$. For this system, the natural dynamics are in general chaotic for sufficiently high levels of initial energy, and the presence of safety-critical effects can be analyzed by inspecting, as before, the positivity of (20). As clear from the plots and predicted by the theory, the controller act purely as a damper (the total energy is non increasing), preserving simple stability of the origin of the phase space. The system converges to solutions with the desired level of energy, which depends both on the choice of $c$ and $\gamma$.
VI Conclusions
In this paper we gave constructive tools to study the qualitative and quantitative effect of safety-critical control schemes implemented with CBFs on the energy balance of controlled physical systems. The analysis led to novel energy-aware schemes, such as selective damping injection mechanisms and active control strategies that inject energy into the controlled system to achieve desired closed-loop behaviors. We are working towards experimental implementations of the proposed algorithms.
[…]